Show that the closed sphere with centre (2,3,7) and radius 10 in 3 R is contained in the
open cube P = {(x, y, z) : x − 2 <11, y − 3 <11, z − 7 <11}.
1
Expert's answer
2020-04-23T17:47:15-0400

We shall show that the closed sphere with centre (2,3,7) and radius 10 in "R^3"  is contained in the

open cube "P = \{(x, y,z): x − 2 <11, y − 3 <11, z − 7 <11\}."

The equation of the sphere with centre (2, 3, 7) and radius 10 in "R^3" is

"(x−2) ^2 +(y−3) ^2 +(z−7) ^2 =10^2\\
(x−2)^2 ≥0,x∈\R\\
(y-3)^2\geq0, y\in \R\\
(z-7)^2\geq0, z\in \R"

Then

"0≤(x−2)^2≤10^2\\
|x-2|\leq10\\
0\leq(y-3)^2\leq 10^2\\
 |y-3|\leq10\\
0\leq(z-7)^2\leq10^2\\
|z-7|\leq10"

Hence

"x-2<11, y-3<11,z-7<11, x,y,z\in R"

This means that the closed sphere with centre (2,3,7) and radius 10 in "R^3" is contained in the

open cube "P = \{(x, y,z): x − 2 <11, y − 3 <11, z − 7 <11\}."